56x^3 + 98x= 170x^2
Answer by mananth(16949)

56x^3 + 98x= 170x^2
56x^3 + 98x- 170x^2=0
2x(28x^2-85x+49)=0
(28x^2-85x+49)=0
using the quadratic equation
a=28, b= -85 , c = 49
The roots x1,x2 = -b +/- sqrt(b^2-4ac) /2a
x1= +85 + sqrt{( -85)^2 - 4*28*49}/2*28
x1= 2.26
.
x2=+85 -sqrt{( -85)^2 - 4*28*49}/2*28
x2= 0.77